### v2

@@ 6 Mitigation Strategies
In this section, we discuss possible mitigation strategies for our attacks and evaluate their impact on the attack success rate as appropriate.
@@ 6.1 Mitigating Query Fingerprinting Attacks
For the query fingerprinting attacks, we propose the following mitigations:
1. Padding Token Packets: Padding the network packet size with additional bytes can mask the actual number of tokens per iteration and conceal the differences between correct and mis-speculations to prevent the attack.
The padding can be added in two ways:
  1. 1.
    Constant Length Padding. Every packet can be padded to a constant size. This size can be selected based on the maximum number of tokens per iteration, multiplied by the maximum bytes per token. We pick this size as 1024 bytes, without loss of generality.
  2. 2.
    Variable Length Padding. Every packet can be padded with a random number of bytes selected from a uniform distribution of $\epsilon\sim Unif(0,D)$ where D is a pre-defined upper bound of padding and $\epsilon$ is an integer.
Table 3: Accuracy after padding each packet to 1024 bytes (attack: 20 traces per prompt; model temperature: 0.8)
No Padding Padding to 1024 Bytes
@@
LADE 13.6% 4%
REST 40% 2%
BiLD 20.4% 2%
Constant Length Padding. Table 3 shows the accuracy of the attack after padding every response packet to a constant size of 1024 Bytes. This completely mitigates the side channel, reducing the attack accuracy to close to random guessing (2%). However, this comes at the cost of considerable communication overhead and inefficiency, due to an increase in packet size by almost 230$\times$, as majority of the packet consists of random padding bytes.
@@
Table 4: Attack accuracy after padding each packets with random bytes from uniform distribution, $\epsilon\sim Unif(0,D)$, where D varies from 6 to 48. (attack: 20 traces per prompt; model temperature: 0.8)
Accuracy vs Max Padding Size (D) Overhead vs D
D $\rightarrow$ 0 6 12 24 48 6 12 24 48
Experiment 1
LADE 65.6% 25.6% 13.6% 8.4% 3.6% 1.5$\times$ 2$\times$ 3.1$\times$ 5.1$\times$
REST 100% 90.4% 75.6% 48% 27.2% 1.9$\times$ 2.8$\times$ 4.6$\times$ 8.2$\times$
BiLD 49.2% 26.8% 11.6% 5.2% 4.8% 1.7$\times$ 2.3$\times$ 3.6$\times$ 6.3$\times$
Experiment 2
LADE 64.8% 19.2% 10.8% 5.6% 4.4% 1.6$\times$ 2.1$\times$ 3.2$\times$ 5.4$\times$
REST 100% 95.2% 76.8% 52% 34.4% 2$\times$ 3$\times$ 4.8$\times$ 8.7$\times$
BiLD 75.2% 37.2% 21.6% 8.4% 5.6% 1.7$\times$ 2.4$\times$ 3.7$\times$ 6.5$\times$
Experiment 3
LADE 13.6% 6.8% 6% 3.6% 3.2% 1.5$\times$ 2$\times$ 3$\times$ 5$\times$
REST 40% 29.6% 20.8% 14.8% 8.4% 1.9$\times$ 2.8$\times$ 4.6$\times$ 8.1$\times$
BiLD 20.4% 7.6% 6.8% 6% 3.6% 1.7$\times$ 2.3$\times$ 3.6$\times$ 6.2$\times$
Variable Length Padding. Table 4 shows the attack accuracy after padding responses with a variable number of bytes, $\epsilon\sim Unif(0,D)$, where D varies from 6 to 48. As D increases to 48, the attack accuracy for LADE and BiLD drops close to random guess (2%) for all experiments. The attack accuracy for REST also shows a considerable drop from 100% without padding to 27% and 34% with padding of up to 48 bytes for Experiments 1 and 2 respectively. Similarly, for Experiment 3, attack accuracy drops from 40% to 8% with D=48. These reductions in accuracy come with a 5$\times$ to 8$\times$ increase in the packet size for D=48. Thus, variable length padding provides a better trade-off of reduction in attack accuracy with limited increase in packet size compared to constant length padding.
@@ D $\rightarrow$ 0 6 12 24 48 6 12 24 48
2. Reduce Granularity of Attacker Observation: An alternative approach is to aggregate output tokens over multiple iterations at the server and return them to the client all together. This can limit observability of fine-grain speculation patterns from the packet sizes for the attacker, and reduce the effectiveness of fingerprinting attacks.
@@
Table 5 shows the results for the server aggregating output tokens over 3 to 20 iterations. The accuracy of the exact knowledge attacks (Experiment 1 and 2) for LADE, REST, and BiLD drops to a minimum of 9.2%, 52%, and 11.2% respectively after aggregating the tokens over up to 20 iterations down from 65%, 100%, and 49.2% respectively without aggregation; accuracy of the approximate knowledge attacks (Experiment 3) decreases to a minimum of 5.2%, 17.6%, and 4% respectively, down from 14%, 40%, and 20% respectively.
Thus, reducing granularity is an effective mitigation against query fingerprinting attacks. One downside of this approach is that the user waits longer between updates to the output as the number of iterations the tokens are aggregated over increases. This limits the number of iterations tokens can be aggregated over without hindering system responsiveness.
@@
Table 5: Accuracy of fingerprinting attacks (20 traces per prompt and temperature of 0.8) with aggregation of tokens across 3 to 20 iterations.
Aggregation Granularity (Number of Iterations)
1 3 5 10 20
Experiment 1
LADE 65.6% 53.2% 40% 24.4% 10.8%
REST 100% 96.8% 89.2% 80.4% 52%
BiLD 49.2% 34% 30% 19.2% 11.2%
Experiment 2
LADE 64.8% 45.2% 28.8% 16.4% 9.2%
REST 100% 97.6% 92.8% 75.6% 52%
BiLD 75.2% 57.6% 41.6% 30% 16.8%
Experiment 3
LADE 13.6% 13.2% 14.4% 9.2% 5.2%
REST 40% 33.2% 28.8% 20.4% 17.6%
BiLD 20.4% 15.6% 10% 10% 4%
6.2 Mitigating Datastore or Hyper-parameter Leaks
For preventing confidential data leakage from datastore or parameter leakage attacks, in addition to the two mitigations discussed previously, we can consider one other mitigation:
3. Using Only Public Data for Speculative Decoding: Any data-store used for speculation (e.g., in REST) should be constructed only with public data, and any private data or personal identifiable information (e.g., usernames, address, etc.) should be scrubbed from it before it is used. This minimizes the risk of sensitive data leakage through any speculation patterns that may be observable despite padding or token aggregation. Similarly, parameters used for speculation (such as N and G in LADE) should be assumed to be public as they can be easily leaked out.
7 Related Work
In this section, we describe other speculative decoding techniques that may be exploitable and related LLM performance optimizations that may lead to side-channel leakage.
7.1 LLM Speculative Decoding Techniques
Speculative Decoding with Tree Attention. SpecInfer Miao et al. (2024) and SpecTr Sun et al. (2023) observe that when a token early in the speculation draft is rejected, the entire draft gets discarded, resulting in wasted computation. Instead, these works propose multiple drafts per iteration organized in a tree structure to provide alternatives for speculative tokens at each position, increasing the likelihood of tokens advancing to the next step and yielding greater speedup by increasing correct speculation rates. These drafts are organized in a tree structure with a custom tree attention mask to prevent inter-tree interactions and verified in parallel. Such works (similar to REST He et al. (2024) which also uses tree attention) increase the correct speculation rate of tokens and introduce more input-dependent speculation patterns in the output. Hence these are likely to be more vulnerable to our query fingerprinting attacks, similar to REST, which is highly susceptible to our attack, as we show in this paper.
Speculative Decoding using Target Models. Other works reuse the target model itself for speculative decoding by utilizing the feature layer (second-to-top layer). Medusa Cai et al. (2024) trains multiple extra heads, each responsible for a token position in the draft sequence, to generate speculative tokens. EAGLE and EAGLE-2 Li et al. (2024a; b) propose a method to perform auto-regressive decoding on the feature layers itself to generate speculative tokens. These techniques are conceptually no different from using an alternative smaller model for speculation like BiLD. Hence, they are equally susceptible to query fingerprinting attacks.
7.2 Performance Optimizations in LLMs
Other performance optimization techniques in LLMs that can introduce timing variation, may also be susceptible to leakage via timing side-channels.
Adaptive KV-Cache Management. The KV-cache is the bottleneck for token generation latency for the memory-bound LLM inference Pope et al. (2023). Recent works, like H2O Zhang et al. (2024b) that performs input-dependent cache eviction or TriForceSun et al. (2024) that performs retrieval-based KV-cache selection for drafting may have input-dependent timing variations. These can introduce timing side-channels allowing query fingerprinting, similar to our attacks. However, as such attacks rely on timing, they are likely to be more susceptible to noise.
Prompt Caching A significant portion of prompts sent to LLMs are duplicates, including system prompts, repeated inputs, and others. Prompt caching, adopted by major LLM providers such as Anthropic and OpenAI Anthropic (2024); OpenAI (2024b); Gim et al. (2024), precomputes attention states for these repetitive prompt components to enable reuse, and substantially reducing latency for the first token when there is reuse. Such techniques can leak the nature of the query based on the timing. Other works cache the responses for prompts with similar semantic meanings Bang (2023). Recent attacks Song et al. (2024) demonstrate query fingerprinting attacks similar to ours, when such prompt caches are shared between users. In contrast, our attacks are effective even without any explicit sharing of caches.
@@ 8 Conclusion
This paper reveals significant privacy risks associated with speculative decoding in Large Language Models. Across multiple speculative decoding techniques, we demonstrate that attacks can exploit speculation patterns to infer user inputs with $>90$% accuracy via fingerprinting attacks and also leak data from datastores used for predictions. These leakages highlight the careful consideration needed when deploying speculative decoding techniques to ensure that performance does not come at the cost of reduced privacy.

### v3

@@ 6.1 Mitigating Query Fingerprinting Attacks
To mitigate the query fingerprinting attacks, we propose (1) padding token packets and (2) aggregating tokens over multiple iterations. We describe both below:
1. Padding Token Packets: Padding the network packet size with additional bytes can mask the actual number of tokens per iteration and conceal the differences between correct and mis-speculations to prevent the attack. The padding can be added in two ways:
[…]
No Padding Fixed Padding Random Padding (bytes) Packet Size Overhead
(1024 bytes) D $\rightarrow$ 6 12 24 48 D $\rightarrow$ 6 12 24 48
@@
[…]
@@
Table 4: Attack accuracy under constant size padding (packets padded to 1024 bytes) and variable-length padding (random padding) mitigations. Variable-length padding follows a uniform distribution, $\epsilon\sim Unif(0,D)$, with $D$ from 6 to 48. (Attack: 20 traces per prompt; model temperature: 0.8).
[…]
@@
2. Reduce Granularity of Attacker Observation: Alternatively, the LLM server can aggregate output tokens over multiple iterations and return them together to the client. This limits observability of fine-grain speculation patterns from packet sizes, reducing the success of fingerprinting attacks.
@@
Table 5 shows the impact of aggregating output tokens over 3 to 20 iterations. The accuracy of the exact knowledge attacks (Experiment 1 and 2) for LADE, REST, and BiLD drops to a minimum of 9.2%, 52%, and 11.2% respectively after aggregating tokens over 20 iterations down from 65%, 100%, and 49.2% respectively without aggregation; accuracy of the approximate knowledge attacks (Experiment 3) decreases to a minimum of 5.2%, 17.6%, and 4% respectively, down from 14%, 40%, and 20%. Thus, token aggregation is an effective mitigation. A downside of aggregating tokens over larger iterations is the increased latency between outputs, which can hinder system responsiveness. At the extreme, aggregating all tokens in a single packet resembles a non-streaming LLM.
Experiment 1 Experiment 2 Experiment 3
AGG $\rightarrow$ 1 3 5 10 20 1 3 5 10 20 1 3 5 10 20
LADE 65.6% 53.2% 40% 24.4% 10.8% 64.8% 45.2% 28.8% 16.4% 9.2% 13.6% 13.2% 14.4% 9.2% 5.2%
REST 100% 96.8% 89.2% 80.4% 52% 100% 97.6% 92.8% 75.6% 52% 40% 33.2% 28.8% 20.4% 17.6%
BiLD 49.2% 34% 30% 19.2% 11.2% 75.2% 57.6% 41.6% 30% 16.8% 20.4% 15.6% 10% 10% 4%
@@
Table 5: Token Aggregation Mitigation. Attack accuracy (20 traces/prompt, temperature of 0.8) as aggregation granularity (AGG), i.e., iterations tokens are aggregated over, varies from 3 to 20.
6.2 Mitigating Datastore Leaks
Use Public Data for Speculation. In addition to padding or token aggregation, to prevent confidential data leakage, a simple approach is to ensure any data-store used for speculation (like in REST) only contains public data, and any private or personal identifiable information (e.g., usernames or address) is anonymized. This eliminates the risk of sensitive data leakage through speculation.
@@ 7 Conclusion
This paper reveals significant privacy risks associated with speculative decoding in Large Language Models. Across multiple speculative decoding techniques, we demonstrate that attacks can exploit speculation patterns to infer user inputs with >90>90% accuracy via fingerprinting attacks on local models and on models remotely hosted on vLLM inference servers, and similarly leak data from datastores used for predictions. These leaks highlight the careful consideration needed when deploying speculative decoding techniques to ensure that performance does not come at the cost of privacy.
@@ A.4 Testing prompts for Experiment 3
A.5 OOD Training prompts for Section 4.8
What are the symptoms of Common cold?
What are the symptoms of Influenza (flu)?
What are the symptoms of COVID-19?
What are the symptoms of Strep throat?
What are the symptoms of Pneumonia?
What are the symptoms of Bronchitis?
What are the symptoms of Tuberculosis (TB)?
What are the symptoms of Urinary tract infection (UTI)?
What are the symptoms of Sexually transmitted infections (chlamydia, gonorrhea)?
What are the symptoms of Lyme disease?
What are the symptoms of Type 2 diabetes?
What are the symptoms of Hypertension (high blood pressure)?
What are the symptoms of High cholesterol?
What are the symptoms of Obesity?
What are the symptoms of Chronic kidney disease?
What are the symptoms of Osteoarthritis?
What are the symptoms of Rheumatoid arthritis?
What are the symptoms of Asthma?
What are the symptoms of Chronic obstructive pulmonary disease (COPD)?
What are the symptoms of Hypothyroidism?
What are the symptoms of Coronary artery disease?
What are the symptoms of Heart failure?
What are the symptoms of Atrial fibrillation?
What are the symptoms of Stroke / transient ischemic attack (TIA)?
What are the symptoms of Peripheral artery disease?
What are the symptoms of Migraine?
What are the symptoms of Epilepsy?
What are the symptoms of Parkinson’s disease?
What are the symptoms of Multiple sclerosis?
What are the symptoms of Alzheimer’s disease / dementia?
What are the symptoms of Gastroesophageal reflux disease (GERD)?
What are the symptoms of Irritable bowel syndrome (IBS)?
What are the symptoms of Peptic ulcer disease?
What are the symptoms of Gallstones?
What are the symptoms of Hepatitis (A, B, C)?
What are the symptoms of Breast cancer?
What are the symptoms of Prostate cancer?
What are the symptoms of Lung cancer?
What are the symptoms of Colorectal cancer?
What are the symptoms of Skin cancer (melanoma, basal cell carcinoma)?
What are the symptoms of Depression?
What are the symptoms of Generalized anxiety disorder?
What are the symptoms of Panic disorder?
What are the symptoms of Bipolar disorder?
What are the symptoms of Post-traumatic stress disorder (PTSD)?
What are the symptoms of Eczema (atopic dermatitis)?
What are the symptoms of Psoriasis?
What are the symptoms of Acne?
What are the symptoms of Rosacea?
What are the symptoms of Fungal skin infections (ringworm, athlete’s foot)?
Appendix B Uniqueness and Reproducibility of Fingerprints
Figure 4 shows the trace of tokens per iteration for two different prompts with LADE (temperature of 0.3) across different runs. By inspecting these t races, we observe two key properties that make these traces suitable for usage to identify and leak the prompt and response:
  1. 1.
    Traces are unique to a prompt and response - different prompts result in different traces of tokens per iteration.
  2. 2.
    Traces are reproducible - repeating the same prompt produces similar patterns in the trace of token counts.
(a) What is narcissistic personality disorder?
(b) What are the symptoms of cancer?
Figure 4: Trace of Tokens per iteration vs Iteration-ID, for two runs of different prompts (a) and (b) with LADE. The trace has unique and reproducible patterns, enabling its usage as a fingerprint for leaking prompts.
Appendix C Leakage of Speculation Hyper-parameters
C.1 Background on LADE’s Implementation
C.2 Leaking Hyper-parameter N in LADE
As each iteration outputs at most N-1 tokens under correct speculation, we expect the maximum number of tokens per iteration to be N-1. Therefore, we prompt LADE with a simple prompt, “Repeat Letter ‘A’ 60 times” that should sustain the maximum correctly speculated tokens per iteration. We observe the maximum number of tokens per iteration with this prompt to deduce N-1, and learn N.
Figure 5(a) shows the number of correct speculated tokens for LADE with three different values of $N$ for this prompt. We see that for each value of N (4, 5, and 6), this prompt has a maximum of 3, 4, and 5 tokens per iteration (N-1), letting us learn the value of N as 4, 5, or 6 in each of these cases. This technique can be extended to learn any value of N used in LADE.
(a) Tokens per iteration with LADE for different NN, with a prompt that achieves maximum correct speculation (“Repeat Letter ‘A’ 60 times”). The maximum tokens per iteration corresponds to N-1.
(b) Tokens per iteration with LADE with Guess Set (G = 3), using a prompt that asks to repeat P phrases sharing a common prefix “I run”. Periodic mis-speculations occur after “run”, only when $P>G$, leaking the value of $G$.
Figure 5: Leaking Speculation Hyper-parameters of LADE.
C.3 Leaking Hyper-parameter G in LADE
To leak G, the number of candidate predictions considered for speculation, we force LADE to incur targeted mis-speculations. Here, we prompt the model to repeat a sequence of phrases with the same prefix (“I run”), i.e., “I run to the grocery shop. I run as fast as cars. I run with my best friend. I run from the lecture hall.”. If the number of such phrases is less than or equal to G, LADE caches all the N-grams following the key “run”, resulting in it being able to correctly speculate the token that follows “run” in each of these phrases. However, if the number of phrases are greater than G, then with LRU replacement, newer n-grams evict older n-grams for the key “run” from the set of candidates which has a limited capacity of G. This causes the token following “run” to always be mis-speculated.
Figure 5(b) shows the traces of token generation time for LADE ($G=3$), as the number of phrases ($P$) of the type “I run …” increases from 2 to 4. When $P$ is 2 or 3 ($\leq G$), we notice all the tokens after the initial ones are correctly predicted in the generation (high number of tokens per iteration). But when $P$ is 4 (>G>G), every 7th token, i.e. the token following “run” is mis-speculated (one token per iteration). Thus, we leak the value of G to be 3. This attack can be generalized to any value of parameter G in LADE.
Appendix D Impact of Mitigation for Token-Length Side Channel
In this experiment, we want to measure the impact on our attacks of mitigation against prior attacks, such as token-length side channels Weiss et al. (2024) that leaks the value of a token based on its associated character length. We evaluate the impact of a mitigation, where every token has been padded to the same size to mask its character length. We measure the attack success rate for our attacks with this mitigation. Note that even with this mitigation, there is variation in the packet sizes, due to simply the number of tokens contained in each packet, which varies due to speculation. Thus, measuring the packet sizes in this scenario is the same as measuring the token counts.
[…]
Temperature
0.3 0.6 0.8 1.0
Experiment 1
LADE 51.6% 38% 27.6% 22.8%
REST 24% 17.2% 10.8% 8.4%
BiLD 67.6% 43.2% 30.4% 22.8%
Experiment 2
LADE 41.2% 24.8% 22.4% 11.2%
REST 25.6% 14% 14.4% 10%
BiLD 82.4% 62.4% 46.8% 33.2%
Experiment 3
LADE 8% 6.8% 6.8% 3.2%
REST 6.4% 4.8% 8% 5.2%
BiLD 8.4% 10% 10.8% 6.8%
Table 6: Accuracy of the Query Fingerprinting Attack with each token padded to the maximum token size (1024 bytes)
Appendix E Out of Distribution Training
Temperature $\rightarrow$ 0.3 0.6 0.8
LADE 25% 24% 23%
REST 36% 35% 36%
BiLD 25% 23% 24%
Table 7: Accuracy of the Query Fingerprinting Attack with out of distribution training, as the temperature of target model varies from 0.3 to 0.8.